# Speeding Up Path Planning via Reinforcement Learning in MCTS for Automated Parking

### v1

@@ I INTRODUCTION
In general, an automated parking task is primarily separated into sub-modules such as perception, localization, planning, control, etc. The appearance of novel perception frameworks such as BEV[1] and occupancy networks[2] brings clarity into sensing the environments in an online automated parking task. There are also well-studied methods of localization to achieve a high-precision result. The performance of an automated parking execution heavily relies on the outcomes from the planning and control module.
@@
Sampling-based approaches are the most common method employed on the path planning task in parking scenarios, and it can be followed by a numeric optimization after. The performance of such an algorithm relies heavily on its sampling density —- the denser the sampling, the algorithm is more likely to converge towards a global optimal solution. Under the parking tasks, especially those complex ones, the vehicle would like to use most of the spare space with bi-directional maneuvers to reach the destination spot with minimal effort while ensuring a comfortable experience for clients. In order to get a satisfying result, a higher dimension of sampling space and a compact sampling strategy are needed, which makes the planning process time-consuming.
Many methods have been introduced to reduce the time consumption for such algorithms in a real-time system. Most of them aim to enhance the exploitation capability of the algorithm, by either trimming out infeasible samples or prioritizing the promising ones with a heuristic estimate. However, too much preference on the exploitation side would decrease the ability to explore and lead to sub-optimal solutions. An accurate estimation during the search steps from prior knowledge would help balance between them.
Monte Carlo tree search (MCTS)[4], as an orderly sampling-based method, shows its strength in solving long-horizon planning tasks like path planning. However, without the significant guidance of prior domain knowledge, the iterative tryouts from its cycled operations make the planning time too long to run in the real-time system. In our approach, we would like to leverage the capability of reinforcement learning to provide domain knowledge guidance for MCTS, boosting up its planning time and finding the equilibrium of searching process, without the help from human driver data.
@@ II RELATED WORKS
The sampling-based approach is well-studied as well. [7] has implemented one of the most popular random-based sampling algorithms: the Rapidly-exploring Random Tree (RRT)[8] for planning on dynamic vehicle model, and it has been used by [9] in automated parking situation. RRT has a simple logic and is easy to implement, but its randomness leads to a lack of stability and is not guaranteed to find an optimal solution.
Besides the random one, another popular orderly-based sampling algorithm is A*[10] and its variant. [11] comes up with a Hybrid-State A* Search which associates a continuous state with each sample, and uses it to solve a perpendicular parking task. The process of A* is expedited by the prior knowledge known as heuristic function[10]. In terms of automated parking, the study on its prior knowledge is mainly focused on the shape properties of the curves, such as the Reeds-Shepp (RS) curve[12], Bezier curve[13], and etc. However, the curve’s shape property becomes less useful when there are lots of obstacles in the parking environment, where Hybrid-State A* works slower compared to the parking task under an open field without obstacles.
Another way to formulate the corresponding prior knowledge considering both the curve’s shape and obstacles is through supervised learning from human expert driver data like [14]. To achieve a well-rounded performance, a large amount of high-quality human data is required and these are often expensive.
@@ III-A Problem Formulation
Equation (1) reveals the composition of action space as an action $a_{t}=(d_{t},\delta_{t})$ , where $\delta_{t}$ is the front wheel angle of the vehicle. (1) itself is the transition function, where $L$ is the length of the wheel base, with the kinematic and collision constraint taken into account.
@@
If the agent at $s_{t}$ reaches the destination pose, $r_{t}=1$, otherwise $r_{t}=0$. The cost function itself is related to safety, comfort, and efficiency of the path formed by traveling through the states.
@@ IV METHODOLOGY
Figure 3: Cycled steps of MCTS in path planning. (a) Iteratively selection tree traversal. (b) Node expansion with policy generator. (c) Simulation with value approximator and cost function. (d) All-way propagation back to root node. (e) Neural network evaluator.
@@ IV-A Monte Carlo Tree Search Design
A tree node $n$ is the combination of a state $s$ and its observation $o$. According to our assumption, with a constant observation $o$, a tree node $n$ has a corresponding state $s$. A tree node $n$ can exist in one of three possible status: UNEXPLORED, EXPLORED, and TRIMMED. A node $n$ is UNEXPLORED when it is just spawned from the expansion of its parent, and it transitions to the EXPLORED status once it is selected and expanded. A node $n$ is TRIMMED if it is trimmed due to certain constraints like collision with obstacles, as the red node shown in Fig. 3, or all children of it are trimmed. A tree node $n$ is considered reaching the destination there is a simple Dubins path[18] can be generated connecting state $s$ and the destination pose.
@@ IV-A3 Simulation
where $\alpha_{0}$ and $\alpha_{1}$ are constant factors which are used to normalize $V$ to $[-1,1]$.
@@ IV-B Reinforcement Learning Design
MCTS itself can be treated as a nature policy improvement operator. Given prior value estimations, it gradually updates its selection policy. After each MCTS is terminated, retrieving the information from the search tree will help us re-estimate the value, as policy evaluation. By iteratively executing policy improvement and policy evaluation process, we have a policy iteration framework where MCTS takes part in both processes.
@@
Figure 4: Reinforcement learning pipeline. Training data is retrieved once MCTS is terminated. The updated $\boldsymbol{p}$ and $v$ are used in the next iteration. The architecture or the evaluation neural network is composed of a convolutional backbone and two separated MLP heads.
@@ V EXPERIMENTS
Derived from real road test scenarios, we constructed thousands of simulated parking scenarios in different types, enumerating the critical features one could have like the size of the parking spots, obstacles’ poses, agent’s starting poses, etc. The generated scenarios are separated into a training dataset and a valid one.
The performance of the MCTS in the inference stage is evaluated based on the planning time it takes to generate a feasible path that meets the path generated by our Hybrid A* algorithm in its quality. We recorded the behavior of Hybrid A* and MCTS at different training steps through the test dataset, as what is shown in Fig. 5. The planning times MCTS takes to find a path drops quickly in the opening stages, and gradually converges to a stable condition. In the end, the median planning time required by MCTS through the dataset is only 7.2% compared to the time taken by Hybrid A*.
@@
Figure 5: Training metric. The darker line is the median over the valid dataset, and the pale shaded area is formed by the 10th and 90th percentiles. The blue one is the planning time which Hybrid A* algorithm takes while the red one is the performance of MCTS at different training steps.
@@ VI CONCLUSION
In this paper, we present a method that integrates reinforcement learning into the Monte Carlo tree search algorithm to expedite automated parking tasks. We formulate path planning as a Markov Decision Process with a bicycle model as the vehicle kinematics. Through the design of the Monte Carlo search tree structure and relative strategies in its iterative steps, we ensure adaptability within the path planning scheme. Furthermore, a neural network is used to generate the policy distribution and estimate the corresponding value of a state within the MCTS. We implement a reinforcement learning pipeline with MCTS serving as a policy improvement operator on our generated dataset, and achieved an excellent training result. The algorithm has been successfully deployed on real-time autonomous vehicles and greatly enhances the experience of the automated parking products.

### v2

@@ I INTRODUCTION
[…]
In general, an automated parking task is primarily separated into sub-modules such as perception, localization, planning, control, etc. The appearance of novel perception frameworks such as Bird’s-Eye-View[1] and occupancy networks[2] brings clarity into sensing the environments in an online automated parking task. There are also well-studied methods of localization to achieve a high-precision result. The performance of an automated parking execution heavily relies on the outcomes from the planning and control module.
@@
Sampling-based approaches are the most common method employed on the path planning task in parking scenarios, and it can be followed by a numeric optimization after. The performance of such an algorithm relies heavily on its sampling density — the denser the sampling, the algorithm is more likely to converge towards a global optimal solution. Under the parking tasks, especially those complex ones, the vehicle would like to use most of the spare space with bi-directional maneuvers to reach the destination spot with minimal effort while ensuring a comfortable experience for clients. In order to get a satisfying result, a higher dimension of sampling space and a compact sampling strategy are needed, which makes the planning process time-consuming.
Many methods have been introduced to reduce the time consumption for such algorithms in a real-time system. Most of them aim to enhance the exploitation capability of the algorithm, by either trimming out infeasible samples or prioritizing the promising ones with a heuristic estimate. However, too much preference on the exploitation side would decrease the ability to explore and lead to sub-optimal solutions.
Monte Carlo tree search (MCTS)[4], as an orderly sampling-based method, shows its strength in solving long-horizon planning tasks like path planning. However, without the significant guidance of prior domain knowledge, the iterative tryouts from its cycled operations make the planning time too long to run in the real-time system. In our approach, we would like to leverage the capability of reinforcement learning to provide domain knowledge guidance for MCTS, boosting up its planning time and finding the equilibrium of searching process, without the help from human driver data, as what is shown in Fig. 1.
@@ II RELATED WORKS
The sampling-based approach is well-studied as well. In [7], implementation of one of the most popular random-based sampling algorithms is given: the Rapidly-exploring Random Tree (RRT)[8] for planning on dynamic vehicle model, and it has been used by [9] in automated parking situation. RRT has a simple logic and is easy to implement, but its randomness leads to a lack of stability and is not guaranteed to find an optimal solution.
Besides the random one, another popular orderly-based sampling algorithm is A*[10] and its variant. [11] comes up with a Hybrid-State A* Search which associates a continuous state with each sample, and uses it to solve a perpendicular parking task. The process of A* is expedited by the prior knowledge known as heuristic function[10]. In terms of automated parking, the study on its prior knowledge is mainly focused on the shape properties of the curves, such as the Reeds-Shepp (RS) curve[12], Bezier curve[13], etc. However, the curve’s shape property becomes less useful when there are lots of obstacles in the parking environment.
Another way to formulate the corresponding prior knowledge considering both the curve’s shape and obstacles is through supervised learning from human expert driver data like [14]. To achieve a well-rounded performance, a large amount of high-quality human data is required and these are often expensive. Prior knowledge can also be enhanced using reinforcement learning strategies. Since the advent of AlphaGo [15], researchers have been working to solve long-horizon planning tasks in the field of autonomous driving, such as those discussed in [16]. [17] managed to address autonomous parking tasks with MCTS, but their solution was limited to parallel parking scenarios.
@@ III-A Problem Formulation
Equation (1) reveals the composition of action space as an action $a_{t}=(d_{t},\delta_{t})$ , where $\delta_{t}$ is the front wheel angle of the vehicle. (1) itself is the transition function, where $L$ is the length of the wheel base, with the kinematic and collision constraint taken into account. In practice, the action space is discretised, which made the state space discrete as well.
@@
If the agent at $s_{t}$ reaches the destination pose, $r_{t}=1$, otherwise $r_{t}=0$. The cost function itself is related to safety, comfort, and efficiency of the path formed by traveling through the states. Paths getting closer to the obstacles within a certain threshold would result in more negative safety cost. Paths with redundant wheel or gear moves would result in more negative comfort and efficiency cost .
@@ IV METHODOLOGY
Figure 3: Cycled steps of MCTS in path planning. (a) Iteratively selection tree traversal. (b) Node expansion with policy generator. (c) Simulation with value approximator and cost function $\mathcal{C}_{\text{path}}$. (d) All-way propagation back to root node. (e) Neural network evaluator.
@@ IV-A Monte Carlo Tree Search Design
A tree node $n$ is the combination of a state $s$ and its observation $o$. According to our assumption, with a constant observation $o$, a tree node $n$ has a corresponding state $s$. A tree node $n$ can exist in one of three possible status: UNEXPLORED, EXPLORED, and TRIMMED. A node $n$ is UNEXPLORED when it is just spawned from the expansion of its parent, and it transitions to the EXPLORED status once it is selected and expanded. A node $n$ is TRIMMED if it is trimmed due to certain constraints like collision with obstacles, as the red node shown in Fig. 3, or all children of it are trimmed. A tree node $n$ is considered reaching the destination if there is a simple Dubins path[20] can be generated connecting its corresponding pose and the corresponding pose of the destination node.
@@ IV-A3 Simulation
where $\alpha_{0}$ and $\alpha_{1}$ are constant factors which are used to normalize $V$ to $[-1,1]$, and $\mathcal{C}_{\text{path}}$ is the cost function.
@@ IV-B Reinforcement Learning Design
MCTS itself can be treated as a natural policy improvement operator. Given prior value estimations, it gradually updates its selection policy. After each MCTS is terminated, retrieving the information from the search tree will help us re-estimate the value, as policy evaluation. By iteratively executing policy improvement and policy evaluation process, we have a policy iteration framework where MCTS takes part in both processes.
@@
Figure 4: Reinforcement learning pipeline. (a) Training data is retrieved once MCTS is terminated, as the policy improvement finished. (b) The architecture of the evaluation neural network is composed of a convolutional backbone and two separated MLP heads. The updated model parameters are used in the next MCTS iteration.
@@ V EXPERIMENTS
Derived from real road test scenarios, we constructed thousands of simulated parking scenarios in different types, enumerating the critical features one could have like the size of the parking spots, obstacles’ poses, agent’s starting poses, etc. The generated scenarios are separated into training, test and validation datasets with a ratio of 70%:15%:15%.
The performance of the MCTS in the inference stage is evaluated based on the planning time it takes to generate a feasible path that meets the path generated by the Hybrid A* [11] algorithm in its quality. We recorded the behavior of Hybrid A* and MCTS at different training steps through the validation dataset, as what is shown in Fig. 5 (a). The discretization setup is 0.1 m for the position and 0.01 Rad for the orientation. The planning time MCTS takes to find a path drops quickly in the opening stages, and gradually converges to a stable condition. In the end, the median planning time required by MCTS through the dataset is only 7.2% compared to the time taken by Hybrid A*. The performance of trained MCTS and Hybrid A* under multiple discretizations is shown in Fig. 5 (b), which suggests that model can be easily scaled up to different setups.
@@
Refer to captionRefer to caption(a)(b)
Figure 5: (a) Training metric. The darker line is the median over the validation dataset, and the pale shaded area is formed by the 10th and 90th percentiles. The blue one is the planning time which Hybrid A* algorithm takes while the red one is the performance of MCTS at different training steps. (b) Planning time and success rate comparison between trained MCTS and Hybrid A* under different discretization setups over the validation dataset.
@@ VI CONCLUSION
In this paper, we present a method that integrates reinforcement learning into the Monte Carlo tree search algorithm to expedite automated parking tasks. We formulate path planning as a Markov Decision Process with a bicycle model as the vehicle kinematics. Through the design of the Monte Carlo search tree structure and relative strategies in its iterative steps, we ensure adaptability within the path planning scheme. Furthermore, a neural network is used to generate the policy distribution and estimate the corresponding value of a state within the MCTS. We implement a reinforcement learning pipeline with MCTS serving as a policy improvement operator on our generated dataset, and achieved an excellent training result. The algorithm has been successfully deployed on real-time autonomous vehicles and greatly enhances the experience of the automated parking product.